OpenStudy (anonymous):

Given f(x) = 4x+1/3, solve for f-1(3).

OpenStudy (anonymous):

2 4 6 8

OpenStudy (dls):

Let f(x)=y \[\LARGE y=4x + \frac{1}{3}\] Solve for x. \[\LARGE x=\frac{3y-1}{4}\] Interchange x and y. So your \[\Large f^{-1} =\frac{3x-1}{4}\] Now,for \(\LARGE f^{-1}(3)=\frac{3(3)-1 }{4}\)=?

OpenStudy (anonymous):

thank you so much! its 2